# Computable Bases

### v1

@@ 4. Closure properties
    $\Diamond:{\mathcal{O}}(X)\to{\mathcal{O}}{\mathcal{A}}_{+}(X),U\mapsto\{A\in{\mathcal{A}}_{+}(X):A\cap U\not$$\emptyset\}$.
@@
We still need to prove that $\mathcal{F}$} and $\Box$ are even computable prebases. Given ${\mathcal{K}}\in{\mathcal{K}}_{-}{\mathcal{K}}_{-}(X)$ we can compute ${\mathrm{sat}}\bigcup_{K\in{\mathcal{K}}}K\in{\mathcal{K}}_{-}(X)$ according to Proposition 17 and because $\bigcap_{K\in{\mathcal{K}}}{\mathcal{F}}_{K}={\mathcal{F}}_{{\mathrm{sat}}\bigcup_{K\in{\mathcal{K}}}K}$ it follows by Lemma 18 that $\mathcal{F}$} is a computable prebase. Likewise, given ${\mathcal{K}}\in{\mathcal{K}}_{-}{\mathcal{O}}(X)$ we can compute $\bigcap_{U\in{\mathcal{K}}}U\in{\mathcal{O}}(X)$ and because $\bigcap_{U\in{\mathcal{K}}}\Box(U)=\Box(\bigcap_{U\in{\mathcal{K}}}U)$ it follows by Lemma 18 that $\Box$ is a computable prebase. ∎
@@
These examples show that in the general case the concept of a computable presubbase and of a computable prebase might actually be more fruitful and interesting than the concept of a computable base. We have lots of natural presubbases and prebases that reveal useful information about our spaces. The presubbase $\mathcal{U}$} does in general not satisfy the convergent intersection property, while the other two presubbases mentioned in Example 27 do.
With the following result we provide a few natural constructions of prebases, where we exploit that the index spaces are overt. We recall that a represented space $X$ is called overt, if $X$ is a computable point in ${\mathcal{A}}_{+}(X)$. Overt spaces are exactly those spaces that allow computable projections on open sets.
@@
(4) Given $U\in{\mathcal{O}}(R\sqcup S)$, where $R\sqcup S=(\{0\}\times R)\cup(\{1\}\times S)$, we can compute the sections $U_{0}\in{\mathcal{O}}(R)$ and $U_{1}\in{\mathcal{O}}(S)$. If $U$ is in the image of $B_{X\sqcup Y}^{\mathsf{{T}}}$ exactly one of $U_{0}$ and $U_{1}$ is non-empty and since $R$ and $S$ are overt, we can find out which one it is. If, for instance, $U_{0}\not$$\emptyset$, then $U_{0}=B_{X}^{\mathsf{{T}}}(x)$ and since $B_{X}^{\mathsf{{T}}}$ is a computable embedding, we can reconstruct $x\in X$ from $U_{0}$. Likewise, we can reconstruct $y\in Y$ in the case of $U_{1}\not$$\emptyset$.
@@
(1) and (2) We prove the claim for $B_{Y^{\mathbb{N}}}$. Given $K\in{\mathcal{K}}_{-}(S^{*})$ we can compute the projection $K_{0}:={\mathrm{pr}}_{0}(K)\in{\mathcal{K}}_{-}({\mathbb{N}})$ onto the natural number component and we can compute some upper bound $m\in{\mathbb{N}}$ on $\max(K_{0})$. Hence, we can compute also the projections $K_{i}:={\mathrm{sat}}({\mathrm{pr}}_{i}(K))\in{\mathcal{K}}_{-}(S)$ on the components i=1,,mi=1,...,m of $K$ by Proposition 17 (which holds uniformly in $i$). For each i=1,,mi=1,...,m we can compute some set $A_{i}\in{\mathcal{A}}_{+}(S)$ with $\bigcap_{s\in K_{i}}B_{Y}(s)=\bigcup_{s\in A_{i}}B_{Y}(s)$. By Proposition 17 (which holds analogously for finite products) we can hence compute $A:=A_{1}\times...\times A_{m}\in{\mathcal{A}}_{+}(S^{*})$ and we obtain for all $n\in{\mathbb{N}}$
@@
where $A_{0}\in{\mathcal{A}}_{+}(R)$ and $A_{1}\in{\mathcal{A}}_{+}(S)$ are sets that we can compute from $K_{0}$ and $K_{1}$ by assumption. Hence we can also compute $A:=A_{0}\sqcup A_{1}\in{\mathcal{A}}_{+}(R\sqcup S)$ because $A\cap U\not$$\emptyset$ for $U\in{\mathcal{O}}(R\sqcup S)$ if and only if $A_{0}\cap U_{0}\not$$\emptyset$ or $A_{1}\cap U_{1}\not$$\emptyset$.
@@
The category of computable Kolmogorov spaces is cartesian closed.
For Corollaries 31 and 32 it was sufficient to have computable presubbases for the respective spaces. The fact that we have even computable prebases for some of these spaces allows further conclusions regarding the underlying topologies. Using Corollaries 16 and 30 and Theorem 25 we obtain the following conclusions. By ${\mathcal{O}}(X)\otimes{\mathcal{O}}(Y)$ and $\bigotimes_{i\in{\mathbb{N}}}{\mathcal{O}}(X)$ we denote the respective product topologies, by ${\mathcal{O}}(X)|_{Z}$ the subspace topology of $Z\subseteq X$ and by ${\mathcal{O}}(X)\wedge{\mathcal{O}}(Y)$ the meet topology generated by the base $B_{X\sqcap Y}$.
@@
Let $X$ and $Y$ be represented ${\mathsf{{T}}}_{0}$ spaces and let $Z\subseteq X$ be a subspace. Then:
@@
We note that for computable Kolmogorov spaces $X$ and $Y$ this result follows from Corollary 30 in combination with Corollary 16. With the help of Theorem 25 we can transfer this result then to arbitrary represented ${\mathsf{{T}}}_{0}$ spaces.
@@ 5. Hyperspace and function space topologies
(1) Let $A\in\Diamond U_{\infty}\cap\bigcap_{n=0}^{k}\Diamond U_{n}$. Then, in particular, $A\cap U_{\infty}\not$$\emptyset$ and hence $A\cap U_{n}\not$$\emptyset$ for all $n\geq k$. This proves the claim.
@@ 6. Epilogue
which was already introduced and discussed in [BR25]. Let us fix a set $X$ and let us denote by $\mathsf{REP}_{0}$ the set of representations $\delta:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$ with a ${\mathsf{{T}}}_{0}$ final topology and let us denote by $\mathsf{PRE}_{0}$ the set of presubbases $B:\subseteq{\mathbb{N}}^{\mathbb{N}}\to 2^{X}$ that induce a ${\mathsf{{T}}}_{0}$ topology on XX. [^3] We call $B:\subseteq{\mathbb{N}}^{\mathbb{N}}\to 2^{X}$ a presubbase if $(B_{y})_{y\in D}$ with $D:={\mathrm{dom}}(B)$ is a presubbase in the sense of our definition.
@@
We also obtain that ${\mathrm{range}}(\Delta)$ is exactly the set of representations of $X$ that turn $X$ into a computable Kolmogorov space (up to computable equivalence) and ${\mathrm{range}}(\nabla)$ is exactly the set of computable bases of $\mathsf{qcb}_{0}$ topologies for $X$ (up to computable equivalence).
@@ Acknowledgments
This work was supported by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) – project number 554999067, by the National Research Foundation of South Africa (NRF) – grant number 151597, and by the Alexander-von-Humboldt Foundation. We would like to thank Matthias Schröder for discussions on the content of this article that have helped to streamline the proofs and their presentations.

### v2

@@ 5. Closure properties
    $\Diamond:{\mathcal{O}}(X)\to{\mathcal{O}}{\mathcal{A}}_{+}(X),U\mapsto\{A\in{\mathcal{A}}_{+}(X):A\cap U\not$$\varnothing\}$.
@@
We still need to prove that $\mathcal{F}$} and $\Box$ are even computable prebases. Given ${\mathcal{K}}\in{\mathcal{K}}_{-}{\mathcal{K}}_{-}(X)$ we can compute ${\mathrm{sat}}\bigcup_{K\in{\mathcal{K}}}K\in{\mathcal{K}}_{-}(X)$ according to Proposition 17 and because $\bigcap_{K\in{\mathcal{K}}}{\mathcal{F}}_{K}={\mathcal{F}}_{{\mathrm{sat}}\bigcup_{K\in{\mathcal{K}}}K}$ it follows by Lemma 18 and Proposition 17 that $\mathcal{F}$} is a computable prebase. Likewise, given ${\mathcal{K}}\in{\mathcal{K}}_{-}{\mathcal{O}}(X)$ we can compute $\bigcap_{U\in{\mathcal{K}}}U\in{\mathcal{O}}(X)$ and because $\bigcap_{U\in{\mathcal{K}}}\Box(U)=\Box(\bigcap_{U\in{\mathcal{K}}}U)$ it follows by Lemma 18 that $\Box$ is a computable prebase. ∎
@@
These examples show that in the general case the concept of a computable presubbase and of a computable prebase might actually be more fruitful and interesting than the concept of a computable base. We have lots of natural presubbases and prebases that reveal useful information about our spaces.
With the following result we provide a few natural constructions of prebases, where we exploit that the index spaces are overt. Overt spaces are exactly those spaces that allow computable projections on open sets.
@@
(4) Given $U\in{\mathcal{O}}(R\sqcup S)$, where $R\sqcup S=(\{0\}\times R)\cup(\{1\}\times S)$, we can compute the sections $U_{0}\in{\mathcal{O}}(R)$ and $U_{1}\in{\mathcal{O}}(S)$. If $U$ is in the image of $B_{X\sqcup Y}^{\mathsf{{T}}}$ exactly one of $U_{0}$ and $U_{1}$ is non-empty and since $R$ and $S$ are overt, we can find out which one it is. If, for instance, $U_{0}\not$$\varnothing$, then $U_{0}=B_{X}^{\mathsf{{T}}}(x)$ and since $B_{X}^{\mathsf{{T}}}$ is a computable embedding, we can reconstruct $x\in X$ from $U_{0}$. Likewise, we can reconstruct $y\in Y$ in the case of $U_{1}\not$$\varnothing$.
@@
(1) and (2) We prove the claim for $B_{Y^{\mathbb{N}}}$. Given $K\in{\mathcal{K}}_{-}(S^{*})$ we can compute the projection $K_{0}:={\mathrm{pr}}_{0}(K)\in{\mathcal{K}}_{-}({\mathbb{N}})$ onto the natural number component and we can compute some upper bound $m\in{\mathbb{N}}$ on $\max(K_{0})$. Hence, we can compute also the projections $K_{i}:={\mathrm{sat}}({\mathrm{pr}}_{i}(K))\in{\mathcal{K}}_{-}(S)$ on the components i=1,,mi=1,...,m of $K$ by Proposition 17 (which holds uniformly in $i$). For each i=1,,mi=1,...,m we can compute some set $A_{i}\in{\mathcal{A}}_{+}(S)$ with $\bigcap_{s\in K_{i}}B_{Y}(s)=\bigcup_{s\in A_{i}}B_{Y}(s)$. By Proposition 17 (which holds analogously for finite products) we can hence compute $A:=A_{1}\times...\times A_{m}\in{\mathcal{A}}_{+}(S^{*})$ and we obtain (with implicitly bound $n\in{\mathbb{N}}$):
@@
where $A_{0}\in{\mathcal{A}}_{+}(R)$ and $A_{1}\in{\mathcal{A}}_{+}(S)$ are sets that we can compute from $K_{0}$ and $K_{1}$ by assumption. Hence we can also compute $A:=A_{0}\sqcup A_{1}\in{\mathcal{A}}_{+}(R\sqcup S)$ because $A\cap U\not$$\varnothing$ for $U\in{\mathcal{O}}(R\sqcup S)$ if and only if $A_{0}\cap U_{0}\not$$\varnothing$ or $A_{1}\cap U_{1}\not$$\varnothing$.
@@
The category of computable Kolmogorov spaces with computable maps as morphisms is cartesian closed.
For Corollaries 48 and 49 it was sufficient to have computable presubbases for the respective spaces. The fact that we have even computable prebases for some of these spaces allows further conclusions regarding the underlying topologies. Using Corollaries 16 and 47 we obtain the following conclusions. By ${\mathcal{O}}(X)\otimes{\mathcal{O}}(Y)$ and $\bigotimes_{i\in{\mathbb{N}}}{\mathcal{O}}(X)$ we denote the respective product topologies, by ${\mathcal{O}}(X)|_{Z}$ the subspace topology of $Z\subseteq X$ and by ${\mathcal{O}}(X)\wedge{\mathcal{O}}(Y)$ the meet topology generated by the base $B_{X\sqcap Y}$.
@@
Let $X$ and $Y$ be admissibly represented ${\mathsf{{T}}}_{0}$ spaces and let $Z\subseteq X$ be a subspace. Then:
@@
This result follows from Corollary 47 in combination with Corollary 16. With the help of Theorem 22 we can transfer (1) and (2) of this result then to arbitrary represented ${\mathsf{{T}}}_{0}$ spaces, since $\delta\mapsto\delta^{\bullet}$ commutes with products [Sch02b].
@@ 6. Hyperspace and function space topologies
(1) Let $A\in\Diamond U_{\infty}\cap\bigcap_{n=0}^{k}\Diamond U_{n}$. Then, in particular, $A\cap U_{\infty}\not$$\varnothing$ and hence $A\cap U_{n}\not$$\varnothing$ for all $n\geq k$. This proves the claim.
@@ 7. Epilogue
Let us fix a set $X$ and let us denote by $\mathsf{REP}_{0}$ the set of representations $\delta:\subseteq{\mathbb{N}}^{\mathbb{N}}\to X$ with a ${\mathsf{{T}}}_{0}$ final topology and let us denote by $\mathsf{PRE}_{0}$ the set of presubbases $B:\subseteq{\mathbb{N}}^{\mathbb{N}}\to 2^{X}$ that induce a ${\mathsf{{T}}}_{0}$ topology on XX. [^3] We call $B:\subseteq{\mathbb{N}}^{\mathbb{N}}\to 2^{X}$ a presubbase if $(B_{y})_{y\in D}$ with $D:={\mathrm{dom}}(B)$ is a presubbase in the sense of our definition.
@@
We also obtain that ${\mathrm{range}}(\Delta)$ is exactly the set of representations of $X$ that turn $X$ into a computable Kolmogorov space (up to computable equivalence) and ${\mathrm{range}}(\nabla)$ is exactly the set of representations of $\mathsf{qcb}_{0}$ topologies for $X$ (up to computable equivalence).
@@ Acknowledgments
This work was supported by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) – project number 554999067, by the National Research Foundation of South Africa (NRF) – grant number 151597, by the Alexander-von-Humboldt Foundation and by the Research Institute for Mathematical Sciences, an international joint research center based at Kyoto University. We would like to thank Matthias Schröder for discussions on the content of this article that have helped to streamline some proofs and their presentations.